OpenStudy (anonymous):

limx->-1 x^3-2x-1/x^5-2x-1 ? without using l'hospital rule

OpenStudy (anonymous):

what do you get when you replace x by -1?

OpenStudy (anonymous):

if you get \[\frac{0}{0}\] then you can factor and cancel. it will be \[\frac{(x+1)\times \text{something}}{(x+1)\times \text{something else}}\]

OpenStudy (anonymous):

cancel and then replace x by -1. in fact it will be \[\frac{x^2-x-1}{x^4-x^3+x^2-x-1}\]

OpenStudy (anonymous):

i got it :) thanks